# Improved Margin Generalization Bounds for Voting Classifiers

### v1

@@ 6 Majority of Three Large Margin Classifiers
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(L_{{\mathbf{S}}_{1}},L_{{\mathbf{S}}_{2}},L_{{\mathbf{S}}_{3}}))\right]$
$\leq\sum_{i>j}\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{i},{\mathbf{S}}_{j}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{i}}({\mathbf{x}}))\not=t({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{j}}({\mathbf{x}}))\not=t({\mathbf{x}})\right]\right]=\frac{288Cd}{\gamma^{2}m}$
@@
We now prove Lemma 12. To the end of showing Lemma 12 we need the following lemma which bounds the conditional error of a large margin learning algorithm under ${\mathcal{D}}_{t}|R$.
@@
There exists a universal constant $c\geq 1$ such that: For distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $t$, hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ with VC-dimension $d$, training sequence size $m$, margin $10<\gamma<$ 1, i.i.d. training sequence ${\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}$, subset $R\subseteq{\mathcal{X}}$ such that $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R\right]:=\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R\right]\not=0$, it holds for any $\gamma$-margin learning algorithm $L\in\operatorname{\Delta({\mathcal{H}})}$ for $t$ that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})\right]\leq\frac{28cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R\right]\gamma^{2}m}.$
@@
For $i\in\{0,1,\ldots\}$ we define the disjoint regions $R_{i}\subseteq{\mathcal{X}}$ $R_{i}=\{x\in{\mathcal{X}}:2^{-i-1}<\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}(x))\not=t(x)\right]\leq 2^{-i}\}$ of ${\mathcal{X}}$ (we will write $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R_{i}\right]$ for short for $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R_{i}\right]$). Then by the law of total probability ${\mathbf{S}}_{1}$ and ${\mathbf{S}}_{2}$ being independent we have
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2}\sim{\mathcal{D}}^{m}_{t}}\negmedspace\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{1}}({\mathbf{x}}))\not=t({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{2}}({\mathbf{x}}))\not=t({\mathbf{x}})\right]\right]\negmedspace=\negmedspace\negmedspace\negmedspace\operatorname*{\mathbb{E}}_{{\mathbf{x}}\sim{\mathcal{D}}}\negmedspace\left[\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}({\mathbf{x}}))\not=t({\mathbf{x}})\right]^{2}\right]$
=$\sum_{i=0}^{\infty}\operatorname*{\mathbb{E}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}({\mathbf{x}})\not=t({\mathbf{x}}))\right]^{2}\Big{|}R_{i}\right]\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R_{i}\right]\leq\sum_{i=0}^{\infty}2^{-2i}\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R_{i}\right],$ (53)
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{1}}({\mathbf{x}}))\not=t({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{2}}({\mathbf{x}}))\not=t({\mathbf{x}})\right]\right]\leq\sum_{i=0}^{\infty}2^{-2i}\frac{8C(i+1)^{2}2^{i}d}{\gamma^{2}m}=\frac{96Cd}{\gamma^{2}m},$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R_{i}}(L_{{\mathbf{S}}})\right]=\operatorname*{\mathbb{E}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}({\mathbf{x}}))\not=t({\mathbf{x}})\right]\Big{|}{\mathbf{x}}\in R_{i}\right]\geq 2^{-i-1},$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R_{i}}(L_{{\mathbf{S}}})\right]\leq\frac{28cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m},$
@@
where the last inequality follows from $C\geq e^{2}.$ Since the above is decreasing in $C$ we get that for $C=(4\cdot 2016)^{2}$, it holds that
@@
If $\frac{d}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}\geq 1$ then we are done by $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}$ always being at most 1, thus for the remainder of the proof we consider the case $\frac{d}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}<1$.
@@
$\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[{\mathbf{N}}\leq\operatorname*{\mathbb{P}}\left[R\right]m/2\right]\leq\exp{\left(-\operatorname*{\mathbb{P}}\left[R\right]m/8\right)}\leq\frac{8}{\operatorname*{\mathbb{P}}\left[R\right]m},$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})\right]\leq\sum_{i=\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil}^{m}\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[{\mathbf{N}}=i\right]+\frac{8}{\operatorname*{\mathbb{P}}\left[R\right]m}.$ (55)
For each $i\in\{\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil,\ldots,m\},$ we will show that $\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]$ is upper bound by $\frac{20cd\ln^{2}{(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d}))}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}$, where $c\geq 1$ is a universal constant, which implies that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})\right]\leq\frac{20cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}+\frac{8}{\operatorname*{\mathbb{P}}\left[R\right]m}\leq\frac{28cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}$
as claimed. We now show for each $i\in\{\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil,\ldots,m\}$ that $\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]\leq\frac{20cd\ln^{2}{(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})))}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}$.
Let for now $i\in\{\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil,\ldots,m\}$. First since $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})$ is nonnegative we have that its expectation can be calculated in terms of its cumulative distribution function,
@@
We now notice that under the conditional distribution ${\mathbf{N}}=i$, it is the case that the $\gamma$- margin learning algorithm $L_{{\mathbf{S}}}$, contains $i$ labelled examples from ${\mathcal{D}}_{t}|R$. Furthermore, since $L_{{\mathbf{S}}}$ has $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(L_{{\mathbf{S}}})=0$ it also has zero margin-loss on the examples drawn from ${\mathcal{D}}_{t}|R.$ Thus, by invoking Theorem 10, it holds with probability at least $1-\delta$ over ${\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}$ conditioned on ${\mathbf{N}}=i$, that
$\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})\leq\frac{cd\operatorname{\mathrm{Ln}}^{2}\left(\frac{\gamma^{2}i}{d}\right)}{\gamma^{2}i}+\frac{c\ln{\left(\frac{e}{\delta}\right)}}{i}\leq\max\left(\frac{2cd\operatorname{\mathrm{Ln}}^{2}\left(\frac{\gamma^{2}i}{d}\right)}{\gamma^{2}i},\frac{2c\ln{\left(\frac{e}{\delta}\right)}}{i}\right)$ (57)
@@
$\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})>\max\left(\frac{2cd\operatorname{\mathrm{Ln}}^{2}\left(\frac{\gamma^{2}i}{d}\right)}{\gamma^{2}i},x\right)\Big{|}{\mathbf{N}}=i\right]\leq e\cdot\exp(-ix/2c),$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]\leq\frac{10cd\ln^{2}{(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d}))}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m},$

### v2

@@ 6 Majority of Three Large Margin Classifiers
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(L_{{\mathbf{S}}_{1}},L_{{\mathbf{S}}_{2}},L_{{\mathbf{S}}_{3}}))\right]$
$\leq\sum_{i>j}\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{i},{\mathbf{S}}_{j}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{i}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{j}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}})\right]\right]=\frac{288Cd}{\gamma^{2}m}$
@@
We now prove Lemma 12. To the end of showing Lemma 12 we need the following lemma which bounds the conditional error of a large margin learning algorithm under ${\mathcal{D}}_{c^{*}}|R$.
@@
There exists a universal constant $c\geq 1$ such that: For distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $c^{*}$, hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ with VC-dimension $d$, training sequence size $m$, margin $10<\gamma<$ 1, i.i.d. training sequence ${\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}$, subset $R\subseteq{\mathcal{X}}$ such that $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R\right]:=\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R\right]\not=0$, it holds for any $\gamma$-margin learning algorithm $L\in\operatorname{\Delta({\mathcal{H}})}$ for $c^{*}$ that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})\right]\leq\frac{28cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R\right]\gamma^{2}m}.$
@@
For $i\in\{0,1,\ldots\}$ we define the disjoint regions $R_{i}\subseteq{\mathcal{X}}$ $R_{i}=\{x\in{\mathcal{X}}:2^{-i-1}<\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}(x))\not=c^{*}(x)\right]\leq 2^{-i}\}$ of ${\mathcal{X}}$ (we will write $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R_{i}\right]$ for short for $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R_{i}\right]$). Then by the law of total probability ${\mathbf{S}}_{1}$ and ${\mathbf{S}}_{2}$ being independent we have
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2}\sim{\mathcal{D}}^{m}_{c^{*}}}\negmedspace\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{1}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{2}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}})\right]\right]\negmedspace=\negmedspace\negmedspace\negmedspace\operatorname*{\mathbb{E}}_{{\mathbf{x}}\sim{\mathcal{D}}}\negmedspace\left[\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}})\right]^{2}\right]$
=$\sum_{i=0}^{\infty}\operatorname*{\mathbb{E}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}({\mathbf{x}})\not=c^{*}({\mathbf{x}}))\right]^{2}\Big{|}R_{i}\right]\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R_{i}\right]\leq\sum_{i=0}^{\infty}2^{-2i}\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[R_{i}\right],$ (53)
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{1}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{2}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}})\right]\right]\leq\sum_{i=0}^{\infty}2^{-2i}\frac{8C(i+1)^{2}2^{i}d}{\gamma^{2}m}=\frac{96Cd}{\gamma^{2}m},$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R_{i}}(L_{{\mathbf{S}}})\right]=\operatorname*{\mathbb{E}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}})\right]\Big{|}{\mathbf{x}}\in R_{i}\right]\geq 2^{-i-1},$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R_{i}}(L_{{\mathbf{S}}})\right]\leq\frac{28cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m},$
@@
where the last inequality follows from $C\geq e^{2}.$ Since the above is decreasing in $C$ we get that for $C=(4\cdot 2016c)^{2}$, it holds that
@@
If $\frac{d}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}\geq 1$ then we are done by $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}$ always being at most 1, thus for the remainder of the proof we consider the case $\frac{d}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}<1$.
@@
$\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[{\mathbf{N}}\leq\operatorname*{\mathbb{P}}\left[R\right]m/2\right]\leq\exp{\left(-\operatorname*{\mathbb{P}}\left[R\right]m/8\right)}\leq\frac{8}{\operatorname*{\mathbb{P}}\left[R\right]m},$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})\right]\leq\sum_{i=\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil}^{m}\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[{\mathbf{N}}=i\right]+\frac{8}{\operatorname*{\mathbb{P}}\left[R\right]m}.$ (55)
For each $i\in\{\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil,\ldots,m\},$ we will show that $\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]$ is upper bound by $\frac{20cd\ln^{2}{(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d}))}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}$, where $c\geq 1$ is a universal constant(the universal constant named $C$ from Theorem 10), which implies that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})\right]\leq\frac{20cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}+\frac{8}{\operatorname*{\mathbb{P}}\left[R\right]m}\leq\frac{28cd\ln^{2}{\left(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})\right)}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}$
as claimed. We now show for each $i\in\{\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil,\ldots,m\}$ that $\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]\leq\frac{20cd\ln^{2}{(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d})))}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}$.
Let for now $i\in\{\left\lceil\operatorname*{\mathbb{P}}\left[R\right]m/2\right\rceil,\ldots,m\}$. First since $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})$ is nonnegative we have that its expectation can be calculated in terms of its cumulative distribution function,
@@
We now notice that under the conditional distribution ${\mathbf{N}}=i$, it is the case that the $\gamma$- margin learning algorithm $L_{{\mathbf{S}}}$, contains $i$ labelled examples from ${\mathcal{D}}_{c^{*}}|R$. Furthermore, since $L_{{\mathbf{S}}}$ has $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(L_{{\mathbf{S}}})=0$ it also has zero margin-loss on the examples drawn from ${\mathcal{D}}_{c^{*}}|R.$ Thus, by invoking Theorem 10, it holds with probability at least $1-\delta$ over ${\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}$ conditioned on ${\mathbf{N}}=i$, that
$\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})\leq\frac{cd\operatorname{\mathrm{Ln}}^{2}\left(\frac{\gamma^{2}i}{d}\right)}{\gamma^{2}i}+\frac{c\ln{\left(\frac{e}{\delta}\right)}}{i}\leq\max\left(\frac{2cd\operatorname{\mathrm{Ln}}^{2}\left(\frac{\gamma^{2}i}{d}\right)}{\gamma^{2}i},\frac{2c\ln{\left(\frac{e}{\delta}\right)}}{i}\right)$ (57)
@@
$\operatorname*{\mathbb{P}}_{{\mathbf{S}}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})>\max\left(\frac{2cd\operatorname{\mathrm{Ln}}^{2}\left(\frac{\gamma^{2}i}{d}\right)}{\gamma^{2}i},x\right)\Big{|}{\mathbf{N}}=i\right]\leq e\cdot\exp(-ix/2c),$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}|R}(L_{{\mathbf{S}}})|{\mathbf{N}}=i\right]\leq\frac{20cd\ln^{2}{(\max(e^{2},\frac{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m}{2d}))}}{\operatorname*{\mathbb{P}}\left[R\right]\gamma^{2}m},$